If a = v^1/2, then at t =0 and at v = 0 find initial velocity(u) ?
1
Expert's answer
2017-12-25T04:39:26-0500
Solution:
Acceleration: a=dv/dt=v^(1/2)
Then, by integrating: dv*v^(-1/2)=dt→2v^(1/2)=t (v(t=0)=0)
So, v(t)=t^2/4
Initial velocity: v(0)=0 (as its said in question)

Answer: 0